there were 500 workers working in a factory.their mean wage was calculated as 200.later on it was discovered that the wage of two workers were misread as 100 and 20 in place of 80 and 220. find the correct average?
Answer by Cromlix(4381)

500 * 200 = Total wages =100,000
100,000 - (100 + 20) = 99880
99880 + (80 + 220) = 100180
100180/500 = 200.36 (New mean)
Hope this helps.
:-)